# Judging LLMs on a Simplex

### v1

@@ 4.1 Probability model
where $n$ is the number of questions. To remove the influence of self-preference, note that (4.2) filters for $k\neq j$. Although judge scores are likely correlated in practice and (4.2) is an overt simplification of the data generating mechanism, the alternative of modeling the full joint distribution introduces additional risks without significant gain: it increases model complexity, requires additional constancy assumptions, risks model misspecification, and yet does not change the identifiability results in Section 3, which were all based on the marginal distribution of judge-assigned scores. As shown in the Appendix, for binary scoring systems, posterior inference under (4.2) yields consistent estimators for $\theta_{m,k}^{(j)}$, prevalences $\vec{\pi_{k}}$, and candidate rankings, even if the correlation structure is ignored.
4.2 Relaxing the constancy assumption
Our identifiability results rest on constancy assumptions that assume shared $\theta_{m}^{(j)}$ parameters. To explore how sensitive rankings are to this assumption, we introduce random effects (RE). There are two options for injecting RE: we can either allow for perturbations in judge performance (i.e. $\theta_{m,k}^{(j)}$ as perturbations of $\theta_{m}^{(j)}$) or score prevalences (i.e. $\vec{\pi}_{k}^{(j)}$ as perturbations of $\vec{\pi}_{k}$). Because the latter involves fewer parameters and thus a lower risk of misspecification, we opt for the latter:
@@ 4.2 Relaxing the constancy assumption
$\begin{split}Z_{k}&\sim\mathrm{Dirichlet}(\delta)\quad\text{{(candidate-specific random direction)}}\\ R_{j}&\sim\mathrm{Beta}(\omega J,J)\quad\text{{(judge random effect)}}\\ W_{k}&\sim\mathrm{Beta}(\omega K,K)\quad\text{{(candidate random effect)}}\\ \vec{\pi}_{k}^{(j)}&=(1-W_{k}R_{j})\vec{\pi}_{k}+W_{k}R_{j}Z_{k},\end{split}$ (4.3)
where the perturbed prevalences $\vec{\pi}_{k}^{(j)}$ would then replace $\vec{\pi}_{k}$ in (4.2). The magnitude of REs is thus controlled through the hyperparameter $\omega\in[0,\infty)$ ($\omega=0$ implies the constancy assumption holds). We recommend setting $\delta$ to detect directions of interest. For instance, to detect over-inflated scores, we set $\delta$ to have larger weights for higher scores (e.g., $\delta=[1,4,10]$ for a 3-level scoring system).
4.3 Injecting prior knowledge about judge quality
By defining a prior over the judge vertices, we can place soft restrictions on the parameter space to exclude regions in which solutions are unlikely to lie. The identifiability results in Section 3 tell us that prevalence rankings can be improved by injecting prior knowledge about likely slopes between judge vertices. This is ideal, as one generally has prior beliefs about the relative locations of the judge vertices but not their absolute locations. By defining a prior over slopes, we minimize the information injected by the prior to only that needed to rank candidates and nothing more and, thus, minimize the risk of having misaligned/misspecified priors.
We use the prior described in Figure 4.4, in which the probability weight from each node is distributed across its children nodes per the transition weights. By designing the transition matrix such that only edges from nodes $(m_{1},m_{2})$ to $(m_{1}^{\prime},m_{2}^{\prime})$ have nonzero weight if $m_{2}\leq m_{2}^{\prime}$, we ensure that Assumption 3 is satisfied. The parameters of the Dirichlet prior $\vec{\beta}_{(m_{1},m_{2})}$ can then either be set based on prior belief or, in our implementation, be sampled from a hyperprior. For instance, for 3-level scoring, we constructed hyperpriors like $\vec{\beta}_{(1,1)}=[1,1+\rho*\beta_{\max},1]$ where $\rho\sim\mathrm{Beta}(1,1)$, where the hyperparameter $\beta_{\max}$ controls the hyperprior over judge quality. See Appendix for details.
@@ 4.3 Injecting prior knowledge about judge quality
Figure 4.4: Bayesian prior for judge vertices $\theta_{m}$, as described using a transition matrix. The transition weights $\alpha_{(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}$ for edges outgoing from the same node must sum to one and are drawn from a Dirichlet prior with parameters $\vec{\beta}_{(m_{1},m_{2})}$. Here we show an example for 3-levels.
5 Experiments
Our theoretical analysis reveals that ranking identifiability depends fundamentally on epistemic and aleatoric uncertainty. Here we evaluate these findings through Bayesian ranking on a simplex, across diverse benchmarks. We find that considering both sources of uncertainty yields more accurate rankings with better-calibrated uncertainty quantification compared to existing methods.
We evaluate the LLM Judge adjudication frameworks across three types of benchmark datasets:
(i) Verifiable tasks with 2-level scores: GPQA [33] and MMLU Pro [45] contain multiple-choice questions, where answers are graded on two levels (correct or incorrect). Judges evaluate candidate answers for correctness with the option to abstain when uncertain.
(ii) Multi-level human-judged tasks: MTBench [50] evaluates multi-turn conversations while TLDR [42] assesses summarizations. Both datasets are assessed on multi-level Likert scales, by LLMs as well as human experts.
(iii) “Semi-verifiable” tasks: Omni-MATH [13] contains mathematical reasoning problems with reference solutions, though there is no single ground truth answer. LLM judges are asked to evaluate answers on a 3-level scale (correct, partial credit, incorrect).
To evaluate judge adjudication methods, we designed a unified two-stage protocol. First, LLM judges evaluate candidates without access to ground truth, mirroring real-world usage. Second, we generate ground truth scores for each answer by comparing against the correct multiple choice answer on verifiable tasks and obtaining human-assigned scores for human-judged tasks. For semi-verifiable tasks, we rescore each candidate’s answer by providing the LLM judge the provided reference answer.
We compare our Bayesian adjudication framework against established baselines: single judge scoring, averaging scores from multiple judges (i.e., LLM judges as a jury), bootstrap confidence intervals [48], and Bradley-Terry pairwise comparison models [31]. For the Bayesian inference procedure, we report results with and without adjustment for self-scoring to examine the impact of potential self-preference bias on rankings and uncertainty estimates. Bayesian inference was implemented using Stan [37] and is extremely fast (less than a minute on a normal laptop). For GPQA, MMLU Pro, and Omni-MATH, we considered 19 contemporary models as candidates, including Claude, GPT, Gemini, Llama, Mistral, and Qwen variants. Their answers were then judged by Claude 3.5 Haiku and GPT-4o mini. We chose these relatively weaker judges to demonstrate that the Bayesian approach remains robust even with less capable evaluators. For MTBench and TLDR, we utilize their existing candidates and judges. Complete candidate and judge lists are available in the Appendix.
Performance was evaluated in terms of Spearman correlation with the ground truth rankings and coverage rates of 95% credible/confidence intervals (averaged across candidates). Because Section 3 highlighted the importance of judge constancy, all analyses are stratified by task characteristics (e.g., GPQA’s undergraduate/graduate/post graduate question levels) to maintain approximate constancy within strata. Additional experiment and implementation details can be found in the Appendix.
Bayesian ranking estimates achieve higher correlations and substantially better coverage than comparator methods (Table 5.1). This is true across all datasets and even different settings of the Bayesian model, illustrating the robustness of the approach. The jump in correlation was highest in GPQA, which is perhaps unsurprising: GPQA was designed to be a difficult dataset, so judge quality is lower in this dataset and rank uncertainty is higher. This is in contrast to datasets like MTBench, where the candidates are so easy to distinguish that the choice of method is less important. That said, because there was evidence of self-preference in MTBench (see Appendix), the adjustment for self-preference was necessary to get perfect correlation with the truth.
Table 5.1: Performance of judge adjudication methods on benchmark datasets. Bayesian ranking on a simplex is run with and without adjustment for self-scoring (adjustment indicated by adj), judge quality hyperparameter $\beta_{\max}=0$, and RE hyperparameter $\omega$=0 or 1. Bayesian ranking consistently achieves both higher correlation (Corr) with ground truth and substantially better coverage rates (Cov) than comparators. Note that there is no self-scoring in TLDR.
Method GPQA MMLU TLDR MTBench Omni-MATH Corr Cov Corr Cov Corr Cov Corr Cov Corr Cov Bayesian: adj, $\omega$=0 0.818 0.852 0.822 0.931 0.885 0.917 1.000 1.000 0.773 0.561 Bayesian: adj, $\omega$=1 0.810 0.852 0.823 0.938 0.907 0.917 1.000 1.000 0.749 0.719 Bayesian: $\omega$=0 0.760 0.778 0.846 0.898 0.971 1.000 0.749 0.649 Bayesian: $\omega$=1 0.760 0.796 0.844 0.901 0.971 1.000 0.726 0.614 Bootstrap 0.720 0.722 0.834 0.898 0.885 0.729 0.971 1.000 0.758 0.544 Bradley-Terry 0.742 0.630 0.841 0.793 0.902 0.771 0.971 1.000 0.756 0.456 Simple Average 0.720 0.093 0.834 0.168 0.885 0.271 0.971 0.833 0.758 0.158 Single Judge 0.696 0.102 0.808 0.151 0.787 0.285 0.986 0.917 0.735 0.202
To better understand ranking identifiability, we visualized results from the Bayesian framework, which revealed the following:
Rankings are generally quite stable across various relaxations of the constancy assumption and judge quality hyperparameters. We conducted Bayesian inference with RE hyperparameter $\omega$=0, 1, 2, 4, and 8, corresponding to expected RE factor magnitudes (i.e., $\mathbb{E}[W_{k}]$) of 0, 0.5, 0.67, 0.8 and 0.89, respectively. Rankings demonstrated notable stability across datasets, with Omni-MATH being the sole exception (Figure 5.5 top left). Indeed, Omni-MATH’s coverage rates reflects this instability, jumping from 0.56 to 0.72 when increasing $\omega$=0 to 1 and reaching 0.90 at $\omega=8$ (Table 5.1). Thus these visualizations can help identify when ranking estimates may be sensitive to modeling assumptions. Variations in the judge quality hyperparameter $\beta_{\max}$ produced minimal impact, suggesting that the data itself contains sufficient information to constrain plausible judge qualities.
Ranking is more uncertain in datasets that are more difficult for LLMs. The posterior distributions in Figure 5.5 (bottom) demonstrate how ranking uncertainty differs across question difficulty tiers in the GPQA dataset. We observe monotonically increasing uncertainty as difficulty increases. This pattern aligns with previous findings that an LLM’s ability to correctly answer questions from a dataset is strongly connected to its ability to judge responses for that dataset [23].
Ranking is more certain when LLM-assigned scores to candidates differ widely. To visualize candidates and judges on a probability simplex, we collapse the benchmarks with 5-point Likert scale into three levels (see Appendix for mapping) (Figure 5.5 top right). The MTBench simplex shows well-separated candidates forming a clear performance gradient, resulting in ranking estimates that are stable due to low epistemic uncertainty. Conversely, TLDR candidates cluster more densely in different regions of the simplex, leading to much higher ranking uncertainty. These contrasting patterns underscore the need to account for both aleatoric and epistemic uncertainty.
Figure 5.5: Visualizing uncertainty. Top left two: Sensitivity of estimated rankings when varying random effects (RE) and judge quality hyperparameters, by plotting the correlation between the estimated ranking for each hyperparameter and their base values. Top right two: Candidates visualized on the probability simplex, when using Claude 3.5 Haiku as a judge on MTBench (left) and GPT-3.5 Turbo as a judge on TLDR (right). Blue triangles are judge configurations sampled from posterior. Bottom: Posterior distributions for candidate rankings across difficulty levels in GPQA.
6 Discussion
This work leverages a novel geometric framework for understanding when gold-standard rankings can and cannot be recovered when one has access to imperfect LLM judges but no gold-standard labels. By visualizing both judges and candidates as points on a probability simplex, we find there is a “phase change”: rankings are generally identifiable for 2-level scoring systems but non-identifiable for 3+-level scoring systems unless prior knowledge is available. This highlights the key roles of aleatoric and epistemic uncertainty when estimating rankings. We then explore how Bayesian inference can be used to integrate both sources of uncertainty and conduct sensitivity analyses by varying the amount of epistemic uncertainty. As the Bayesian framework treats judges as black boxes, it is compatible with any judge ensemble. Empirically, we find that Bayesian ranking on a simplex consistently outperforms existing baselines in both correlation with and coverage of the true rankings.
This work suggests the following analysis protocol for practitioners using LLM judge rankings:
Recommendations for judge analysis: Score all LLM candidates using two or more LLM judges. Primary analysis: Perform Bayesian posterior inference on all non-self-judged scores assuming zero random effects. Obtain posterior mean and credible intervals for candidate rankings. Constancy sensitivity analysis: Test sensitivity to the constancy assumption by varying RE hyperparameter $\omega$. Plot how much ranking estimates vary across different values of $\omega$. Judge quality sensitivity analysis: Assess the impact of judge quality by varying the judge quality hyperparameter $\beta_{\max}$. Plot how much ranking estimates vary across different values of $\beta_{\max}$.
Future work. The primary goal of this work is to understand the theoretical limits of rank estimation and, thus, considered a relatively simple Bayesian model. Future directions may consider more sophisticated models that, say, incorporate the correlation structure to further improve estimation and uncertainty quantification. Other extensions include extending the Bayesian model to hybrid systems with some gold-standard labels and understanding how this model can help strategically allocate human effort to improve ranking estimates [2, 6].
Impact statement. This work highlights both potential risks and benefits of using LLMs as judges in the absence of human labels. Recommendations in this work can be used alongside benchmarking frameworks and leaderboards such as HELM [27] to improve uncertainty quantification and measure sensitivity to assumptions. Findings in this work can also help LLM practitioners and researchers take a more deliberate approach when leveraging these technologies to scale LLM evaluation.
Appendix
Appendix A Additional figures
Figure A.6: Example reasons why ranking may still be identifiable when there are 3+ levels. Left: Candidates have assigned scores that are very different. Right: If a judge satisfies strong constancy and has judged a lot of candidates, the range of judge performance characteristics becomes very limited.
Appendix B Bayesian model
B.1 Model description
The full Bayesian model is specified by the plate diagram in Figure B.7. As described in Section 4.3 of the main text, Dirichlet prior over the transition weights $\vec{\alpha}_{m}$ is parameterized by $\vec{\beta}_{m}$. For a setting with $3$ true scores, the parameters are defined as follows:
@@ B.1 Model description
where $\beta_{\max}$ is a hyperparameter that controls the prior over judge quality. (Note that there is no need for a hyperprior for $\alpha_{1,3}$ and $\alpha_{2,3}$ as they are both equal to one.) Extensions to fewer or more levels is follows a similar format (see code for exact implementation).
Figure B.7: Plate diagram of the Bayesian model
@@ Appendix C Implementation details
Bayesian inference: We conduct posterior inference using Hamiltonian Monte Carlo (HMC), coded up using Stan [37]. HMC ran with 4 chains, each running 1000 iterations for warmup and another 1000 for sampling.
@@ D.1 Benchmarks
  • GPQA [33]: Questions resistant to simple internet searches across STEM domains, stratified by difficulty (undergraduate, graduate, post-graduate).
  • MMLU Pro [45]: Enhanced professional knowledge questions across 16 domain-specific subcategories from natural sciences, social sciences, and humanities.
  • MTBench [50]: A conversational benchmark evaluating single- and multi-turn dialogue capabilities across diverse scenarios (creative writing, reasoning, coding, mathematics, role-playing). Human judges rated responses on a 10-point scale. For LLM judges, we ask for ratings on a simplified 5-point scale.
  • TLDR (aka SummEval) [11]: A summarization benchmark where models condense news articles into concise summaries, with human ratings across four dimensions (relevance, consistency, fluency, coherence) on 5-point scales.
  • Omni-MATH [13]: A benchmark focusing on high-difficulty competition-level problems from International and National Olympiads. These problems present particular challenges for automated evaluation as solutions vary significantly in approach, notation, and presentation; multiple valid solution paths may exist for the same problem; and partial correctness must be assessed along multiple dimensions.
D.1.1 Strata Definitions for Benchmark Datasets
To maintain approximate judge constancy within evaluation groups, we stratify each benchmark dataset based on task characteristics that are likely to affect judge performance consistency. The stratification approach differs across benchmarks based on the available metadata and inherent structure of each dataset.
GPQA Stratification: For GPQA, we stratify questions according to the difficulty levels assigned by the original question writers. This stratification yields three distinct strata:
  • Hard Undergraduate: Questions designed to challenge advanced undergraduate students in STEM fields
  • Hard Graduate: Questions requiring graduate-level knowledge and reasoning
  • Post Graduate: Questions at the level expected for advanced graduate students or researchers
MMLU Pro Stratification: For MMLU Pro, we focus on a subset of subjects that include educational level metadata, allowing us to stratify by both subject domain and academic level. We consider seven major subject areas, each divided into high school and college levels where available:
  • Physics: High school physics, College physics
  • Mathematics: High school statistics, High school mathematics, College mathematics
  • History: High school US history, High school European history, High school world history
  • Economics: High school macroeconomics, High school microeconomics
  • Computer Science: High school computer science, College computer science
  • Chemistry: High school chemistry, College chemistry
  • Biology: High school biology, College biology
Omni-MATH Stratification: For Omni-MATH, we use the competition tier structure inherent in the benchmark data. Each competition is mapped to one of three tiers (T0, T1, T2) based on the typical difficulty and selectivity of the mathematical olympiad:
  • Tier 0 (T0): Highest difficulty competitions including IMO, IMO Shortlist, IMC, and Putnam
  • Tier 1 (T1): High-level national competitions such as USAMO, national olympiads, and team selection tests
  • Tier 2 (T2): Regional and junior competitions including USAJMO, JBMO, APMO, and Balkan MO

### v2

@@ B.1 Probability Model
Although judge scores are likely correlated in practice and (4) simplifies the true data generating mechanism, modeling the full joint distribution would increase complexity and require additional constancy assumptions without changing the identifiability results in Section 3. As we show in Theorem 5, for binary scoring systems, posterior inference under this model yields consistent estimators for prevalences and rankings even when correlation structure is ignored.
B.2 Random Effects Parameterization
B.2.1 Complete Hierarchical Specification
To relax the constancy assumption while maintaining computational tractability, we introduce random effects that allow judge-specific and candidate-specific deviations from base prevalences. The complete hierarchical model is:
@@ B.2.1 Complete Hierarchical Specification
$Z_{k}$ $\sim\text{Dirichlet}(\delta)\quad\text{(candidate-specific random direction)}$ (5)
$R_{j}$ $\sim\text{Beta}(\omega J,J)\quad\text{(judge-specific random effect magnitude)}$ (6)
$W_{k}$ $\sim\text{Beta}(\omega K,K)\quad\text{(candidate-specific random effect magnitude)}$ (7)
$\vec{\pi}_{k}^{(j)}$ $=(1-W_{k}R_{j})\vec{\pi}_{k}+W_{k}R_{j}Z_{k}\quad\text{(perturbed prevalences)}$ (8)
The perturbed prevalences $\vec{\pi}_{k}^{(j)}$ replace $\vec{\pi}_{k}$ in the likelihood (4). This parameterization ensures that $\vec{\pi}_{k}^{(j)}$ remains on the probability simplex for all valid parameter values.
B.2.2 Choice of Prevalence vs. Confusion Matrix Perturbation
We considered two approaches for introducing random effects: we can either allow for perturbations in judge performance (i.e. $\theta_{m,k}^{(j)}$ as perturbations of $\theta_{m}^{(j)}$) or score prevalences (i.e. $\vec{\pi}_{k}^{(j)}$ as perturbations of $\vec{\pi}_{k}$).
We chose prevalence perturbation for several reasons. For $M$-level scoring with $J$ judges and $K$ candidates, confusion matrix perturbation requires $O(JKM^{2})$ parameters while prevalence perturbation requires only $O(JKM)$. Fewer parameters reduce the risk of overfitting and model misspecification and enables faster convergence in MCMC sampling. Finally, deviations in prevalences directly correspond to changes in perceived candidate quality by different judges.
B.2.3 Setting $\delta$ for Detecting Specific Biases
The Dirichlet parameter $\delta$ controls the distribution of random directions and can be tuned to detect specific types of bias:
  • Uniform exploration: $\delta=[1,1,...,1]$ explores all directions equally
  • Score inflation detection: For 3-level scoring, $\delta=[1,4,10]$ places more weight on higher scores, detecting judges who systematically overrate candidates
  • Score deflation detection: For 3-level scoring, $\delta=[10,4,1]$ places more weight on lower scores, detecting overly critical judges
  • Central tendency bias: $\delta=[1,10,1]$ for 3-level scoring detects judges who avoid extreme scores
In our experiments, we use score inflation detection ($\delta=[1,4,10]$) as self-preference bias typically manifests as inflated scores for one’s own model family.
B.3 Judge Quality Prior Specification
The judge quality prior is implemented through a transition weight framework that parameterizes the confusion matrix columns (judge vertices) as weighted combinations of base positions. This section provides a detailed mathematical specification of the framework briefly introduced in Section 5.3 of the main paper. This approach allows us to encode beliefs about relative vertex positions while maintaining the probabilistic constraints of the confusion matrix.
B.3.1 Transition Weight Framework
We can parameterize the judge’s confusion matrix columns (vertices) through a directed acyclic graph where nodes represent confusion matrix entries and edges control how probability mass flows between them. This allows us to encode prior beliefs about judge quality through the edge weights while ensuring mathematical constraints are satisfied.
@@ B.3.1 Transition Weight Framework
Figure B.10: Transition weight framework for encoding judge quality priors in 3-level scoring. Nodes represent pairs $(m_{1},m_{2})$ where $m_{1}$ is the true score and $m_{2}$ is the assigned score. Edges show allowed transitions, with weights $\alpha$ drawn from Dirichlet priors parameterized by $\vec{\beta}_{(m_{1},m_{2})}$ (see Appendix B.3 for formal specification).
The formal specification of the transition weights is:
$\left\{\alpha_{(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}\right\}_{(m_{1}^{\prime},m_{2}^{\prime})\text{ s.t. }(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}\sim\text{Dirichlet}(\vec{\beta}_{(m_{1},m_{2})})$
In this notation, each node $(m_{1},m_{2})$ in the transition graph represents an entry in the confusion matrix: the probability of assigning score $m_{2}$ when the true score is $m_{1}$. The transition weights $\alpha_{(m_{1},m_{2})\rightarrow(m_{1}^{\prime},m_{2}^{\prime})}$ control how probability mass flows from parent nodes to child nodes. All outgoing weights from any parent node must sum to one, ensuring the result remains a valid probability distribution. The final confusion matrix entry $\theta_{m_{1}^{\prime},m_{2}^{\prime}}^{(j)}$ for judge $j$ is computed as a weighted average of its parent nodes’ values, weighted by the incoming edge weights.
For example, in Figure B.10, the probability $\theta_{2,2}$ (assigning score 2 when true score is 2) receives contributions from both $\theta_{1,1}$ and $\theta_{1,2}$ through their respective edge weights.
B.3.2 Parameterization for 3-Level Scoring
For a 3-level scoring system, the Dirichlet parameters $\vec{\beta}_{(m_{1},m_{2})}$ are specified as follows:
@@ B.3.2 Parameterization for 3-Level Scoring
where:
  • $\beta_{\max}$ is the global hyperparameter controlling prior strength about judge quality
  • $\rho^{(j)}\sim\text{Beta}(1,1)$ allows judge-specific variation in discrimination ability (drawn from a uniform hyperprior on $[0,1]$)
  • The structure ensures higher weight on correct identification (diagonal elements) as $\beta_{\max}$ increases
Note that $\alpha_{1,3}$ and $\alpha_{2,3}$ are deterministically set equal to one as these represent terminal nodes in the transition graph with no outgoing edges.
Intuitively, when $\beta_{\max}=0$, all Dirichlet parameters equal 1 (uniform prior), allowing maximum flexibility in judge behavior. As $\beta_{\max}$ increases, the parameters $1,1+\rho^{(j)}\beta_{\max},1]$ place more weight on the middle transition, encoding our belief that better judges more consistently assign higher scores when the true score increases.
B.3.3 Enforcing Monotonicity Through Graph Structure
The transition graph topology automatically enforces Assumption 3 (monotonicity) by restricting which edges exist. Specifically, edges from node $(m_{1},m_{2})$ to $(m_{1}^{\prime},m_{2}^{\prime})$ only exist when:
  • $m_{1}^{\prime}=m_{1}+1$ (moving to the next true score level)
  • $m_{2}^{\prime}\geq m_{2}$ (assigned score cannot decrease as true score increases)
This structural constraint ensures that judges cannot be “label-flippers” who systematically assign lower scores to better answers, without requiring explicit inequality constraints during MCMC sampling.
B.3.4 Extension to $M$-Level Scoring
For general $M$-level scoring systems, the framework extends naturally:
  • The graph has $M(M+1)/2$ nodes representing all valid $(m_{1},m_{2})$ pairs where $m_{2}\geq m_{1}$
  • Edge constraints follow the same monotonicity rules
  • Dirichlet parameters follow the pattern: higher values along paths that preserve or increase the assigned score relative to the true score
The specific parameterization for $M>3$ follows similar principles, with $\beta_{\max}$ controlling the concentration around correct identification. See the accompanying code for exact implementation details for different values of $M$.
@@ Appendix C Implementation details
Bayesian inference: We conduct posterior inference using Hamiltonian Monte Carlo (HMC), coded up using Stan (Stan_Development_Team2021-id). HMC ran with 4 chains, each running 1000 iterations for warmup and another 1000 for sampling.
@@ D.1 Benchmarks
  • GPQA (Rein2023-dv): Questions resistant to simple internet searches across STEM domains, stratified by difficulty (undergraduate, graduate, post-graduate).
  • MMLU Pro (Wang2024-ic): Enhanced professional knowledge questions across 16 domain-specific subcategories from natural sciences, social sciences, and humanities.
  • MTBench (Zheng2023-wr): A conversational benchmark evaluating single- and multi-turn dialogue capabilities across diverse scenarios (creative writing, reasoning, coding, mathematics, role-playing). Human judges rated responses on a 10-point scale. For LLM judges, we ask for ratings on a simplified 5-point scale.
  • TLDR (aka SummEval) (Fabbri2021-ie): A summarization benchmark where models condense news articles into concise summaries, with human ratings across four dimensions (relevance, consistency, fluency, coherence) on 5-point scales.
  • Omni-MATH (Gao2024-ec): A benchmark focusing on high-difficulty competition-level problems from International and National Olympiads. These problems present particular challenges for automated evaluation as solutions vary significantly in approach, notation, and presentation; multiple valid solution paths may exist for the same problem; and partial correctness must be assessed along multiple dimensions.